# NPHardEval: Dynamic Benchmark on Reasoning Ability of Large Language Models via Complexity Classes

### 25 Dec 2023 v2

@@ 5.3 Effects of Few-shot Examples’ Difficulty on Reasoning Ability Enhancement
Zeroshot 0.000 0.000 0.000 0.000 0.440 0.000 0.040 0.960 0.160 0.950
Fewshot (-5) 0.000 0.000 0.000 0.140 0.160 0.000 0.140 0.880 0.460 0.640
Fewshot (-4) 0.000 0.000 0.000 0.100 0.150 0.000 0.050 0.817 0.383 0.483
Fewshot (-3) 0.000 0.043 0.000 0.057 0.100 0.000 0.029 0.871 0.329 0.400
Fewshot (-2) 0.000 0.038 0.000 0.025 0.075 0.000 0.025 0.700 0.238 0.350
Fewshot (-1) 0.000 0.011 0.000 0.056 0.033 0.000 0.000 0.667 0.200 0.256
Fewshot (0) 0.000 0.020 0.000 0.060 0.000 0.000 0.000 0.730 0.190 0.200
Fewshot (1) 0.000 0.000 0.000 0.070 0.000 0.000 0.000 0.800 0.190 0.210
Fewshot (2) 0.000 0.040 0.000 0.050 0.000 0.000 0.000 0.750 0.090 0.100
Fewshot (3) 0.000 0.020 0.000 0.060 0.040 0.000 0.000 0.710 0.040 0.100
Fewshot (4) 0.000 0.030 0.000 0.030 0.000 0.000 0.000 0.810 0.000 0.000
Fewshot (5) 0.000 0.050 0.000 0.030 0.000 0.000 0.000 0.820 0.030 0.000
@@ 6 Conclusions and Discussion
In this study, we present a novel benchmark, NPHardEval, designed to rigorously evaluate LLMs’ reasoning capabilities across a spectrum of complex tasks, up to the complexity class of NP-hard. By eschewing standard QA formats in favor of complex, logic-oriented problems, this benchmark aims to provide a more accurate measure of a model’s reasoning prowess. This approach is crucial for developing LLMs capable of handling sophisticated, real-world tasks that demand high-level cognitive processing, steering the evaluation of LLMs from potentially ‘useful’ to fundamentally ‘logical’.
@@ 6.2 Research Outlook
Robustness Evaluation
As mentioned in the limitations, a critical area for future research is the robustness testing of LLMs. This involves examining how variations in parameters like ‘temperature’ in GPT models openai2023gpt4 , which controls the randomness of response generation, impact the model’s performance. A low temperature results in more predictable and conservative responses, while a higher temperature encourages diversity and creativity in the responses. Understanding how these settings affect the reasoning abilities of LLMs can provide deeper insights into their operational mechanisms. Additionally, conducting multiple iterations of the same question with the same model is essential to assess the consistency and reliability of the models. This repeated testing can help in identifying patterns and anomalies in the model’s reasoning process, further contributing to the robustness analysis.
Dynamic Update Plan
As we cast our sights to the future, the NPHardEval benchmark is set to incorporate a dynamic update mechanism—a feature critical to its evolution and relevance. This mechanism is meticulously designed to refresh the datapoints every month. The rationale for such regular updates is twofold: firstly, it ensures that LLMs are evaluated on current and varied task sets, thereby reducing the potential for overfitting to static benchmarks. Secondly, it allows for the tracking of LLMs’ performance over time, offering insights into their learning curves and adaptability to new types of tasks.
The dynamic update plan is not merely a maintenance protocol but a strategic initiative to continuously challenge and refine the reasoning faculties of LLMs. By introducing new tasks of varied complexities and contexts, we anticipate that the LLMs will be subjected to a more robust and comprehensive testing environment. This, in turn, should yield a more accurate and reliable assessment of their capabilities.
@@
Moreover, exploring a multi-agent system approach could significantly advance LLMs’ reasoning abilities. In such a system, different LLM agents, each potentially specialized in certain types of reasoning or knowledge areas, collaborate to solve complex problems. This collaborative approach could mimic a team of experts, each contributing their expertise, leading to more comprehensive and nuanced solutions. It also opens the door to understanding how LLMs can interact and augment each other’s capabilities, which is crucial for their application in real-world, multi-faceted problem-solving scenarios.

### 12 Jan 2024 v3

@@ 5.4 Effects of Few-shot Examples’ Difficulty on Reasoning Ability Enhancement
Zeroshot 0.000 0.000 0.000 0.000 0.440 0.000 0.000 0.040 0.000 0.960 0.160 0.950
Fewshot (-5) 0.000 0.000 0.000 0.140 0.160 0.000 0.320 0.140 0.540 0.880 0.460 0.640
Fewshot (-4) 0.000 0.000 0.000 0.100 0.150 0.000 0.233 0.050 0.400 0.817 0.383 0.483
Fewshot (-3) 0.000 0.043 0.000 0.057 0.100 0.000 0.100 0.029 0.300 0.871 0.329 0.400
Fewshot (-2) 0.000 0.038 0.000 0.025 0.075 0.000 0.038 0.025 0.263 0.700 0.238 0.350
Fewshot (-1) 0.000 0.011 0.000 0.056 0.033 0.000 0.022 0.000 0.167 0.667 0.200 0.256
Fewshot (0) 0.000 0.020 0.000 0.060 0.000 0.000 0.040 0.000 0.130 0.730 0.190 0.200
Fewshot (1) 0.000 0.000 0.000 0.070 0.000 0.000 0.010 0.000 0.100 0.800 0.190 0.210
Fewshot (2) 0.000 0.040 0.000 0.050 0.000 0.000 0.020 0.000 0.000 0.750 0.090 0.100
Fewshot (3) 0.000 0.020 0.000 0.060 0.040 0.000 0.000 0.000 0.010 0.710 0.040 0.100
Fewshot (4) 0.000 0.030 0.000 0.030 0.000 0.000 0.000 0.000 0.000 0.810 0.000 0.000
Fewshot (5) 0.000 0.050 0.000 0.030 0.000 0.000 0.000 0.000 0.000 0.820 0.030 0.000
@@ 6 Conclusions and Discussion
In this study, we present a novel benchmark, NPHardEval, designed to rigorously evaluate LLMs’ reasoning capabilities across a spectrum of complex tasks, up to the complexity class of NP-hard. By eschewing standard QA formats in favor of complex, logic-oriented problems, this benchmark aims to provide a more accurate measure of a model’s reasoning prowess. This approach is crucial for developing LLMs capable of handling sophisticated, real-world tasks that demand high-level cognitive processing, steering the evaluation of LLMs from potentially “useful" to fundamentally “logical".
@@ 6.2 Research Outlook
Moreover, exploring a multi-agent system wu2023autogen ; chan2023chateval ; ge2023openagi ; ge2023llm approach could significantly advance LLMs’ reasoning abilities. In such a system, different LLM agents, each potentially specialized in certain types of reasoning or knowledge areas, collaborate to solve complex problems. This collaborative approach could mimic a team of experts, each contributing their expertise, leading to more comprehensive and nuanced solutions. It also opens the door to understanding how LLMs can interact and augment each other’s capabilities, which is crucial for their application in real-world, multi-faceted problem-solving scenarios.
@@
7 Acknowledgement
We extend our sincere gratitude to Libby Hemphill for her invaluable support in this work. Additionally, we are grateful for the diverse online feedback we received, which illuminated our path towards enhancing the quality of this paper.